OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 0) and a radius of 5.

OpenStudy (freckles):

The equation of a circle comes from the distance formula \[r^2=(x-h)^2+(y-k)^2 \] where (h,k) is the center and r is the radius

OpenStudy (anonymous):

so it would be 5^2=(x-0)^2+(y-0)^2

OpenStudy (anonymous):

wait no... x^2+y^2-25=0

OpenStudy (freckles):

yep that is right both ways you wrote it is right

OpenStudy (anonymous):

thank you!

OpenStudy (freckles):

np